A glass of water is filled to a height of 0.500 ft. Find the hydrostatic pressure at the bottom of the glass.


1
Expert's answer
2021-02-02T18:47:57-0500

By the definition of hydrostatic pressure, we get:


"P=\rho gh+P_0,"

"P=1000\ \dfrac{kg}{m^3}\cdot9.8\ \dfrac{m}{s^2}\cdot0.5\ ft\cdot\dfrac{0.3048\ m}{1\ ft}+1.013\cdot10^5\ Pa=1.028\cdot10^5\ Pa."